OpenStudy (anonymous):

Question: Write the result in scientific notation. (1.1 x 10^-5) (3 x 10^-2) A. 4.1 x 10 ^-7 B. 4.1 x 10^10 C. 3.3 x 10^-7 D. 3.3 x 10^10 Optional: If you can, can you help me by explaining how you got this so I can understand? Thank You!

OpenStudy (anonymous):

the answer is D

OpenStudy (anonymous):

(1.1 x 10^-5) (3 x 10^-2) So 10^x * 10^y= 10^(x+y) just do 1.1*3 and the total of 10^-5*10^-2 This means -5+(-2)=-7

OpenStudy (anonymous):

this is because both have the same basis:10

OpenStudy (anonymous):

Wait, so that would mean that its C because -5 + -2 = -7 making it 10 ^ -7.

OpenStudy (anonymous):

Yeap, thats true!!! Sorry for answer you D lol

OpenStudy (anonymous):

Oh is alright, thanks!

OpenStudy (anonymous):

ok